You are researching a company and find that it has an inventory period of 22.4 days, an accounts payable period of 36.5 days, and an accounts receivable period of 31.4 days. The company's cash cycle is ___ days and its operating cycle is ___ days.
Answer #1

Cash cycle is the period which is required for business to convert the raw material into cash.

Cash cycle = Inventory period + accounts receivable period - Accounts payable period

Cash cycle = 22.4 + 31.4 - 36.5

Cash cycle = 17.3 days

Operating cycle is the period required to receive inventory, sale it and collect cash

Operating cycle = inventory period + accounts receivable period

Operating cycle = 22.4 + 31.4 = 53.8 days